4x + 5y = 51
4x - 5y = 21

Answer by Fombitz(32388)

Add the equations together.
4x+5y+4x-5y=51+21
8x=72
x=9
THen use either equation to solve for y.
4x+5y=51
4(9)+5y=51
36+5y=51
5y=15
y=3
(x,y)=(9,3)